93447 is 1000 more than?
66000 is 100 more than?
13206 is 1000 less than?
4998 is 100 less than?
1
Expert's answer
2013-02-28T11:17:50-0500
93447 is 1000 more than?
93447-1000=92447
93447 is 1000 more than 92447

66000 is 100 more than?
66000-100=65900
66000 is 100 more than 65900

13206 is 1000 less than?
13206+1000=14206
13206 is 1000 less than 14206

4998 is 100 less than?
4998+100=5098
4998 is 100 less than 5098